### v1

# Learning-guided Kansa collocation for forward and inverse PDEs beyond linearity

Abstract
Partial Differential Equations are precise in modelling the physical, biological and graphical phenomena. However, the numerical methods suffer from the curse of dimensionality, high computation costs and domain-specific discretization. We aim to explore pros and cons of different PDE solvers, and apply them to specific scientific simulation problems, including forwarding solution, inverse problems and equations discovery. In particular, we extend the recent Zhong et al. (2023) framework solver to multi-dependent-variable and non-linear settings, together with down-stream applications. The outcomes include implementation of selected methods, self-tuning techniques, evaluation on benchmark problems and a comprehensive survey of neural PDE solvers and scientific simulation applications.
1 Introduction
PDEs are useful in different domains of scientific computing, including physics, graphics and biology. They describe how quantities change over space and time, making them essential for modeling phenomena such as fluid dynamics, electromagnetic fields, heat transfer, and population dynamics.
Zhong et al. (2023) proposed extension to Kansa method, which is a mesh-free Radial Basis Functions (RBFs) PDE solver. They introduced auto-tuning of the shape parameters of RBFs. However, their work focuses only on single-variable linear PDEs. Therefore, this paper extends CNFs backend solver to multi-dependent-variable $u$ and nonlinear PDEs, and apply the framework to specific scientific simulation problems, including forward computation and inverse problems.
It’s unknown how (extended) CNF solvers Zhong et al. (2023) compared with other classical and neural PDE solvers. Hence, we also implement and evaluate selected prior methods on the benchmarks on their effectiveness with different quality metrics (e.g. L1, L2, errors) against ground truth solutions, efficiency, computation resource, convergence speed, method complexity, and finally utility in research, i.e. their scientific simulation applications or integration with other methods, e.g. differentiable rendering to solve inverse physics-related problems in Graphics.
2 Related work
PDE benchmarks. We identified several representative equations Takamoto et al. (2022) in Table 14. They are different in linearity of the operator and solution $u$ dimensionality.
PDE solvers. Numerical methods, e.g. Finite Difference Method (FDM) and Finite Element Method (FEM), are widely used to solve PDEs. However, they suffer from the curse of dimensionality, high computation costs and domain-specific discretization. Recently, neural network based solvers have shown promising results in addressing these issues. For example, Physics-Informed Neural Networks (PINNs) Raissi et al. (2019) and Fourier Neural Operators (FNOs) Li et al. (2020) have demonstrated the ability to generalize to unseen scenarios and handle high dimension effectively.
Inverse problem, i.e. estimating unknown parameters or inputs of a variable $x$ from given solution observations $u$, is crucial. However, it’s unclear how CNFs can be applied to these problems, including connecting with differentiable rendering pipelines Spielberg et al. (2023) in Visual Computing.
3 Methodology
3.1 General form of PDEs
With spatial domain $\Omega\subset\mathbb{R}^{d}$, where its dimension is $d$, and the unknown field $u(x,t)\in\mathcal{U}:\mathbb{D}\to\mathbb{R}$ defined on the spatio-temporal domain $\mathbb{D}=\Omega\times[t_{0},t_{f}]\subset\mathbb{R}^{d+1}$, the general form of PDEs is,
$\begin{cases}\begin{aligned} \mathcal{D}[u]=f,&\quad x\in\Omega,t\in[t_{0},t_{f}],\\ \mathcal{B}_{i}[u]=g_{i},&\quad x\in\partial\Omega_{i},t\in[t_{0},t_{f}].\end{aligned}\end{cases}\Leftrightarrow\begin{cases}\begin{aligned} \mathcal{D}[u](x,t)=f(x,t),&\quad x\in\Omega,t\in[t_{0},t_{f}],\\ \mathcal{B}_{i}[u](x,t)=g_{i}(x,t)&\quad x\in\partial\Omega_{i},t\in[t_{0},t_{f}].\end{aligned}.\end{cases}\\$ (1)
where $\mathcal{D}:\mathcal{U}\to\mathcal{Y}$ is the differential operator and $f\in\mathcal{Y}:\mathbb{D}\to\mathbb{R}^{m}$ is source function, e.g. the external force in dynamics, with $m$ being the output dimension of ff [^1] Note that $\mathcal{U}$ and $\mathcal{Y}$ are two function spaces, and we require they are Banach spaces.. The differential operators $\mathcal{D}$ include the gradient $\nabla$, Laplace $\Delta$, divergence $\nabla\cdot$, etc. For boundary conditions, $\mathcal{B}_{i}:\mathcal{U}\to\mathcal{Z}_{i}$ is each boundary operator with $g_{i}\in\mathcal{Z}_{i}:\partial\Omega_{i}\times[t_{0},t_{f}]\to\mathbb{R}^{n_{i}}$ and $n_{i}$ as the output dimension of $g_{i}$.
3.2 Kansa collocation
Kernel functions. Radial Basis Function (RBF) inversely relates the distance $r$ between the input $\mathbf{x}$ and a fixed origin point $\mathbf{c}$ to the output value.
[…]
There are various infinitely smooth RBFs, among which we choose Gaussian RBF for its effectiveness in approximating smooth functions,
$\psi_{\mathbf{c}}(r)=\begin{cases}e^{-(\epsilon r)^{2}},&\quad\text{Gaussian},\\ \frac{1}{1+(\epsilon r)^{2}},&\quad\text{Inverse quadratic},\\ \sqrt{1+(\epsilon r)^{2}},&\quad\text{Multiquadrics},\end{cases}$ (3)
where Gaussian shape parameter $\epsilon=\frac{1}{\sqrt{2}\sigma}$, and $\sigma$ is the standard deviation.
[…]
where $\alpha_{i}\in\mathbb{R}$ are the coefficients to be solved. The time dimension $t$ is omitted, which can be treated as an additional spatial dimension here. Equation equation 4 is expressed as, by rewriting the kernel functions into matrix form,
[…]
The PDE general form equation 1 can be summarized as a single equation,
$\mathcal{F}[\hat{u}](\mathbf{x}_{i})=h(\mathbf{x}_{i}),\quad\mathbf{x}_{i}\in\mathbb{D},$ (6)
where the operator $\mathcal{F}=\{\mathcal{D},\mathcal{B}_{i}\}$ and $h=\{f,g_{i}\}$ represent both the initial and boundary conditions.
3.2.1 Linear operator case
By plugging in the approximation of $u$ equation 4 and assuming the operator $\mathcal{F}$ is linear [^2] For linear operators, $\mathcal{F}[\alpha\cdot\psi]=\alpha\cdot\mathcal{F}[\psi]$, as defined in § A., the PDE equation 6 can be simplified as,
$\mathcal{F}[\hat{u}](\mathbf{x}_{i})=\mathcal{F}[\sum_{i=1}^{N}\alpha_{i}\cdot\psi_{i}](\mathbf{x}_{i})=\sum_{i=1}^{N}\alpha_{i}\cdot\mathcal{F}[\psi_{i}](\mathbf{x}_{i})=h(\mathbf{x}_{i}).$ (7)
By expanding in matrix form, the above equation is,
$\underbrace{\begin{bmatrix}\mathcal{F}[\psi_{1}](\mathbf{x}_{1})&\mathcal{F}[\psi_{2}](\mathbf{x}_{1})&\cdots&\mathcal{F}[\psi_{N}](\mathbf{x}_{1})\\ \mathcal{F}[\psi_{1}](\mathbf{x}_{2})&\mathcal{F}[\psi_{2}](\mathbf{x}_{2})&\cdots&\mathcal{F}[\psi_{N}](\mathbf{x}_{2})\\ \vdots&\vdots&\ddots&\vdots\\ \mathcal{F}[\psi_{1}](\mathbf{x}_{N})&\mathcal{F}[\psi_{2}](\mathbf{x}_{N})&\cdots&\mathcal{F}[\psi_{N}](\mathbf{x}_{N})\end{bmatrix}}_{\text{operator-evaluated kernel matrix }\mathbf{F}\in\mathbb{R}^{N\times N}}\cdot\underbrace{\begin{bmatrix}\alpha_{1}\\ \alpha_{2}\\ \vdots\\ \alpha_{N}\end{bmatrix}}_{\mathbf{a}\in\mathbb{R}^{N}}=\underbrace{\begin{bmatrix}h(\mathbf{x}_{1})\\ h(\mathbf{x}_{2})\\ \vdots\\ h(\mathbf{x}_{N})\end{bmatrix}}_{\text{constraint values }\mathbf{h}\in\mathbb{R}^{N}}.$ (8)
Concretely, when the kernel function is Gaussian RBF defined in equation 2, and $\mathbf{x}_{i}=(x_{i},t_{i})$,
$\psi_{i}=e^{-\frac{r_{i}^{2}}{2\sigma^{2}}},\quad r_{i}^{2}=(x-x_{i})^{2}+(t-t_{i})^{2}.$ (9)
[…]
$\mathcal{F}[\psi_{i}](\mathbf{x}_{j})=\frac{\partial\psi_{i}(\mathbf{x}_{j})}{\partial t}=\frac{\partial\psi_{i}(\mathbf{x}_{j})}{\partial r_{i}^{2}}\cdot\frac{\partial r_{i}^{2}}{\partial t}=-\frac{1}{2\sigma^{2}}e^{-\frac{r_{i}^{2}}{2\sigma^{2}}}\cdot 2(t_{j}-t_{i})=-\frac{t_{j}-t_{i}}{\sigma^{2}}e^{-\frac{r_{i}^{2}}{2\sigma^{2}}}.$ (10)
Simultaneous equations. When there are $N_{eq}$ equations of different constraints to be satisfied, the collocation points $\{\mathbf{x}_{i}\in\mathbb{D}\}_{i=1}^{N_{\text{total}}}$ are distributed among all equations [^4] Note that repeated collocation points are forced to be repeated here for distinct constraints., where $N_{\text{total}}=\sum_{j=1}^{N_{eq}}N_{j}$.
$\mathcal{F}_{j}[\hat{u}](\mathbf{x}_{i})=h_{j}(\mathbf{x}_{i}),\quad\forall j\in\{1,\ldots,N_{eq}\},\mathbf{x}_{i}\in\mathbb{D}.$ (11)
Equation equation 8 can be extended by stacking each matrix $\mathbf{F}^{(j)}\in\mathbb{R}^{N_{\text{total}}\times N_{\text{total}}}$ and constraint vector $\mathbf{h}^{(j)}\in\mathbb{R}^{N_{\text{total}}}$ vertically for all equations Zhong et al. (2023). The block matrix form is,
$\underbrace{\begin{bmatrix}\mathbf{F}^{(1)}\\ \vdots\\ \mathbf{F}^{(N_{eq})}\end{bmatrix}}_{\text{stacked }\mathbf{F}\in\mathbb{R}^{(N_{eq}\cdot N_{\text{total}})\times N_{\text{total}}}}\cdot\underbrace{\begin{bmatrix}\alpha_{1}\\ \vdots\\ \alpha_{N_{\text{total}}}\end{bmatrix}}_{\mathbf{a}\in\mathbb{R}^{N_{\text{total}}}}=\underbrace{\begin{bmatrix}\mathbf{h}^{(1)}\\ \vdots\\ \mathbf{h}^{(N_{eq})}\end{bmatrix}}_{\text{stacked }\mathbf{h}\in\mathbb{R}^{N_{eq}\cdot N_{\text{total}}}}.$ (12)
The solution of $u$ equation 4 depends on the coefficients $\mathbf{a}=[\alpha_{1},\alpha_{2},\ldots,\alpha_{N}]$, which can be solved by the linear system $\mathbf{F}\mathbf{a}=\mathbf{h}$. The general form is given by the least squares approximation, i.e. minimizing the norm of the error vector and setting the gradient to zero,
$\mathbf{a}^{\text{opt}}$ $=\min_{\mathbf{a}}(||\mathbf{F}\mathbf{a}-\mathbf{h}||)^{2},$ (13)
$\nabla_{\mathbf{a}}(\mathbf{F}\mathbf{a}-\mathbf{h})^{T}(\mathbf{F}\mathbf{a}-\mathbf{h})=0\implies(\mathbf{F}^{T}\mathbf{F})\mathbf{a}^{\text{opt}}=\mathbf{F}^{T}\mathbf{h}.$
If matrix $\mathbf{F}$ is full rank, $\mathbf{F}^{T}\mathbf{F}$ is invertible, thus one can derive $\mathbf{a}^{\text{opt}}=(\mathbf{F}^{T}\mathbf{F})^{-1}\mathbf{F}^{T}\mathbf{h}$. Should the matrix $\mathbf{F}$ be square and invertible, equation 13 can be further simplified as $\mathbf{a}^{\text{opt}}=\mathbf{F}^{-1}\mathbf{h}$. Whichever conditions occurs, the final solution for $u$ is approximated by plugging in the optimal coefficients $\mathbf{a}^{\text{opt}}$ into equation 5, i.e. $\hat{u}(\mathbf{x})=\mathbf{K}\cdot\mathbf{a}^{\text{opt}}$.
When testing on unseen data points $\{\mathbf{x}_{j}^{\star}\in\mathbb{D}\}_{j=1}^{M}$, the kernel functions are constructed between the test points and the collocation points $\{\mathbf{x}_{i}\in\mathbb{D}\}_{i=1}^{N}$. The solution is thus,
[…]
Test-time solution equation 14 can be formulated to matrix form,
[…]
3.2.2 Extension 1: coupled solution fields of PDEs
Coupled multi-dimensional PDE solution fields. Assuming there are $N_{D}$ solution dimensions, i.e. $\mathbf{u}=[u_{1},u_{2},\ldots,u_{N_{D}}]$, the Kansa approximation equation 4 for each dimension is,
[…]
The couple PDE equation is thus formulated as applying the coupling, or governing operator $\mathcal{G}$ on all dimensions of solution, which each has its own operator $\mathcal{F}^{(d)}$,
$\mathcal{G}\left(\mathcal{F}^{(1)}[\hat{u}_{1}],\ldots,\mathcal{F}^{(N_{D})}[\hat{u}_{N_{D}}]\right)(\mathbf{x}_{i})=h(\mathbf{x}_{i}),\quad\mathbf{x}_{i}\in\mathbb{D}.$ (17)
Here we assume the coupling operator $\mathcal{G}$ is linear with each dimension of solution,
$\mathcal{G}\left(\hat{v}_{1},\ldots,\hat{v}_{N_{D}}\right)(\mathbf{x})=\sum_{d=1}^{N_{D}}\beta_{d}\cdot\hat{v}_{d}(\mathbf{x}),$ (18)
where $\beta_{d}\in\mathbb{R}$ is the per-dimension weight. Equation equation 8 can be extended by stacking each matrix $\mathbf{F}^{(d)}\in\mathbb{R}^{N\times N}$ horizontally for all dimensions of solution. The block matrix form is,
$\underbrace{\begin{bmatrix}\beta_{1}\textbf{I}_{N}&\cdots&\beta_{N_{D}}\textbf{I}_{N}\end{bmatrix}}_{\bm{\beta}\in\mathbb{R}^{N\times(N_{D}\cdot N)}}\circ\underbrace{\begin{bmatrix}\mathbf{F}^{(1)}&\cdots&\mathbf{F}^{(N_{D})}\end{bmatrix}}_{\text{coupling }\mathbf{F}\in\mathbb{R}^{N\times(N_{D}\cdot N)}}\cdot\underbrace{\begin{bmatrix}\textbf{a}^{(1)}\\ \vdots\\ \textbf{a}^{(N_{D})}\end{bmatrix}}_{\mathbf{a}\in\mathbb{R}^{(N_{D}\cdot N)}}=\underbrace{\begin{bmatrix}h(\mathbf{x}_{1})\\ \vdots\\ h(\mathbf{x}_{N})\end{bmatrix}}_{\text{stacked }\mathbf{h}\in\mathbb{R}^{N}},$ (19)
where $\circ$ is element-wise or Hadamard product and $\textbf{I}_{N}$ is the identity matrix of size $N$. For simultaneous coupled PDE equations, similar to equation 11, they are indexed by $j\in\{1,\ldots,N_{eq}\}$,
$\mathcal{G}_{j}\left(\mathcal{F}_{j}^{(1)}[\hat{u}_{1}],\ldots,\mathcal{F}_{j}^{(N_{D})}[\hat{u}_{N_{D}}]\right)(\mathbf{x}_{i})=h_{j}(\mathbf{x}_{i}),\quad\forall j\in\{1,\ldots,N_{eq}\},\mathbf{x}_{i}\in\mathbb{D}.$ (20)
With $N_{\text{total}}$ defined as in equation 12, the block matrix form is,
$\underbrace{\begin{bmatrix}\bm{\beta}^{(1)}\\ \vdots\\ \bm{\beta}^{(N_{eq})}\end{bmatrix}}_{\bm{\beta}\in\mathbb{R}^{(N_{eq}\cdot N_{\text{total}})\times(N_{D}\cdot N_{\text{total}})}}\circ\underbrace{\begin{bmatrix}\mathbf{F}^{(1,1)}&\cdots&\mathbf{F}^{(1,N_{D})}\\ \vdots&\mathbf{F}^{(j,d)}&\vdots\\ \mathbf{F}^{(N_{eq},1)}&\cdots&\mathbf{F}^{(N_{eq},N_{D})}\end{bmatrix}}_{\text{coupling }\mathbf{F}\in\mathbb{R}^{(N_{eq}\cdot N_{\text{total}})\times(N_{D}\cdot N_{\text{total}})}}\cdot\underbrace{\begin{bmatrix}\textbf{a}^{(1)}\\ \vdots\\ \textbf{a}^{(N_{D})}\end{bmatrix}}_{\mathbf{a}\in\mathbb{R}^{(N_{D}\cdot N_{\text{total}})}}=\underbrace{\begin{bmatrix}\textbf{h}^{(1)}\\ \vdots\\ \textbf{h}^{(N_{eq})}\end{bmatrix}}_{\text{stacked }\mathbf{h}\in\mathbb{R}^{(N_{eq}\cdot N_{\text{total}})}}.$ (21)
3.2.3 Extension 2: nonlinear operator case
When the operator $\mathcal{F}$ is nonlinear, we can no longer simplify equation 6 as in equation 7. However, we can still derive the relation between the solution $u$ and its linear transformed version as below.
Differentiable matrix helps decompose the general non-linear operator $\mathcal{F}$ into a series of linear operators. Take any linear operator, e.g. $\frac{\partial}{\partial x}$, it relates the relation between unknown u and its derivative $\mathbf{u}^{\prime}=\mathbf{D}_{x}\cdot\mathbf{u}$. We derive $\mathbf{D}_{x}$ from Kansa equation 4, by linearity and equation 7,
[…]
We can invert equation 5 $\mathbf{a}=\mathbf{K}^{-1}\cdot\mathbf{u}$, assuming $\mathbf{K}$ invertibility from independent basis. By substituting $\mathbf{a}$ into $\mathbf{u}^{\prime}=\mathbf{K_{x}}\cdot\mathbf{a}$, one gets $\mathbf{u}^{\prime}=\mathbf{K_{x}}\cdot\mathbf{K}^{-1}\cdot\mathbf{u}$. The differentiable matrix is thus,
$\mathbf{D}_{x}=\mathbf{K_{x}}\cdot\mathbf{K}^{-1}\in\mathbb{R}^{N\times N}.$ (23)
For viscous Burgers’ equation equation 63 $\mathcal{F}[u]=\frac{\partial u}{\partial t}+u\frac{\partial u}{\partial x}-\nu\frac{\partial^{2}u}{\partial x^{2}}$. Its differentiable matrix form, which follows the same formulation as in equation 23 by replacing the operator accordingly, is,
$\mathcal{F}[u]=\mathbf{D}_{t}\cdot\mathbf{u}+\mathbf{u}\circ(\mathbf{D}_{x}\cdot\mathbf{u})-\nu(\mathbf{D}_{xx}\cdot\mathbf{u}).$ (24)
Here we present two categories of Kansa approaches (Table LABEL:tab:nonlinear-kansa-error-bound). The first consists of four time-stepping schemes, including two per-step linear and another two nonlinear systems. The second employs a fully nonlinear solver on the PDE residuals, without explicit time discretization.
Table 1: Summary of different non-linear Kansa solver features, $\Delta t$ is the time step size, $N_{x}$ and $N_{t}$ are the number of collocation points in spatial and temporal dimensions respectively.
Features forward IMEX backward Crank–Nicolson fully non-linear
Time-step explicit semi-explicit implicit implicit $\times$
Error $O(\Delta t)$ $O(\Delta t)$ $O(\Delta t)$ $O(\Delta t^{2})$ $O(1)$
Stability unstable stable stable stable N/A
Memory $O(N_{x}^{2})$ $O(N_{x}^{2})$ $O(N_{x}^{2})$ $O(N_{x}^{2})$ $O(N_{x}^{2}N_{t}^{2})$
Time-stepping approach with linear system.
We can remove the non-linearity by discretizing the time derivative via finite difference method, for a special case of time-dependent PDEs. One solution is to use the (1) explicit forward Euler scheme,
$\frac{\partial u}{\partial t}+\mathcal{D}[u]=0\implies\frac{u^{n+1}-u^{n}}{\Delta t}+O(\Delta t)+\mathcal{D}[u^{n}]=0,$ (25)
where $\mathcal{D}$ is the spatial operator. A more stable solution is to use the (2) implicit-explicit (IMEX) scheme, which splits the stiff and non-stiff parts of the operator $\mathcal{D}=\mathcal{I}_{\text{stiff}}+\mathcal{E}_{\text{non-stiff}}$, which stiffness means the numerical instability incurred by the operator and needs to be treated implicitly,
$\frac{u^{n+1}-u^{n}}{\Delta t}+O(\Delta t)+\mathcal{I}_{\text{stiff}}[u^{n+1}]+\mathcal{E}_{\text{non-stiff}}[u^{n}]=0.$ (26)
Despite the non-linear spatial operator $\mathcal{D}$ or $\mathcal{E}_{\text{non-stiff}}$, we already know the solution $u^{n}$ at time step $n$. Thus, with differentiable matrices, one can evaluate $\mathcal{D}[u^{n}]$ or $\mathcal{E}_{\text{non-stiff}}[u^{n}]$ directly, so as to derive the solution $u^{n+1}$ at the next time step $n+1$.
Time-stepping approach with nonlinear solver. If we discretize the time derivative via the (3) backward Euler scheme, the non-linearity remains in the formulation,
$\frac{\partial u}{\partial t}+\mathcal{D}[u]=0\implies\frac{u^{n+1}-u^{n}}{\Delta t}+O(\Delta t)+\mathcal{D}[u^{n+1}]=0.$ (27)
We can directly replace linear system solver by a non-linear system solver, e.g. Newton-Raphson method Ypma (1995), to minimize the residual vector and derive the unknown solution at next time step,
[…]
Alternatively, (4) Crank-Nicolson scheme can be used to discretize second-order accurate in time,
$\frac{\partial u}{\partial t}+\mathcal{D}[u]=0\implies\frac{u^{n+1}-u^{n}}{\Delta t}+\frac{1}{2}\left(\mathcal{D}[u^{n+1}]+\mathcal{D}[u^{n}]\right)+O(\Delta t^{2})=0.$ (29)
Similar with equation 28, the unknown solution at next time step is derived by minimizing the residual vector as stated in equation 29.
Fully nonlinear solver without time-stepping. This approach directly minimizes the PDE residuals equation 6 over all collocation points, without explicit time discretization. After plugging in the differentiable matrix form of the non-linear operator $\mathcal{F}$, the objective function is therefore,
$\alpha=\arg\min_{\alpha}\sum_{i=1}^{N}\left(\mathcal{F}[\hat{u}](\mathbf{x}_{i})-h(\mathbf{x}_{i})\right)^{2}.$ (30)
By plugging in Kansa approximation equation 4, we derive unknown solution $u$ over entire domain.
3.2.4 Auto-tuning of Kansa hyperparameters
To tune the key Kansa method hyperparameter, kernel shape parameter $\epsilon$ in equation 3, Zhong et al. (2023) proposed one of the self-tuning methods for $\epsilon$, by minimizing the variation of the solution field $u$ over all collocation points, and the condition number of the kernel matrix $\mathbf{F}$,
[…]
where $\text{cond}(\mathbf{F})$ is the condition number of matrix $\mathbf{F}$ defined in equation 8. The integral term can be approximated by summing over all collocation points by Monte Carlo integration. This approach works for linear, including coupled and multi-dimensional, PDEs.
For non-linear operator case, the solution $u$ depends on $\epsilon$ implicitly via the coefficients $\alpha_{i}$. The matrix $\mathbf{F}$ no longer exists explicitly. Here, we propose to directly minimize the PDE residuals over all collocation points, the total variation of the solution field $u$, and the training L2 loss between the predicted solution $u$ and the ground truth solution $u^{gt}$ if available,
[…]
where $\omega_{1},\omega_{2}$, and $\omega_{3}$ are the weights for each penalty term.
3.3 Solutions of inverse PDE problems
Inverse PDE problems.
When given observations of solution field $u^{\text{obs}}$, we infer the unknown PDE parameters $\bm{\pi}$ that minimize the discrepancy $\mathcal{L}$ between the predicted $u^{\text{pred}}(\bm{\pi})$ and $u^{\text{obs}}$,
$\bm{\pi}^{\star}=\arg\min_{\bm{\pi}}\mathcal{L}(u^{\text{obs}},u^{\text{pred}}(\bm{\pi})).$ (33)
We adopt the SciPy implementation of the least squares and root finding algorithms, which are either gradient-based or gradient-free, detailed in the evaluation section.
4 Evaluation
4.1 Performance metrics
Accuracy. Given the numerical solution $\hat{u}_{i}$ from PDE solvers, and the corresponding ground truth $u_{i}$, the $L_{2}$ risk $\mathcal{R}_{L_{2}}$ is the average discretized error over all $N_{\text{test}}$ test points,
[…]
The relative $L_{2}$ risk is computed from $\mathcal{R}_{L_{2}}$ and normalized by the ground truth $u_{i}$ $L_{2}$ norm,
4.2 Evaluation of solvers for the Advection equation
For the 1D advection equation defined in equation 45, we set the number of domain quadrature points $N_{\mathcal{R}}=100\times 10$, i.e. initial condition (IC) points $N_{d}=10$ and the boundary condition (BC) points $N_{\mathcal{B}}=100\times 2$. The advection equation is initialized as per Table LABEL:tab:advection-1d-setup.
Table 2: 1D advection equation experimental setup.
domain time range parameter IC BC
$x_{0}=0,x_{f}=1$ $t_{0}=0,t_{f}=1$ $\beta=0.4$ $u_{0}(x)=\sin(2\pi x)$ per equation 46
FNO requires multiple instances of PDEs for training. Hence, we generate $N_{pde}=100$ instances by varying only the initial condition as, given $c_{k}\sim\mathcal{N}(0,1)$,
$u_{0}(x):=\frac{u_{0}(x)}{\max_{x}|u_{0}(x)|},\text{ where }u_{0}(x)=\sum_{k=1}^{5}c_{k}\,\sin(2\pi kx).$ (35)
[…]
4.2.1 Forward problem
Since FNO is trained on $N_{pde}=100$ instances of PDEs, we compensate more training data for single-instance solvers for a fair comparison. The adjustment factor is defined as $C_{\text{scale}}\in[1,N_{pde}]\subset\mathbb{R}^{+}$. Hence, the domain points is $N^{\prime}_{\mathcal{R}}=C_{\text{scale}}\times N_{\mathcal{R}}$ and methods denoted as $\text{FDM}^{C_{\text{scale}}}$ and $\text{PINN}^{C_{\text{scale}}}$. The test-time results are summarized in Table LABEL:tab:advection-1d-error.
[…]
$C_{\text{scale}}$ FDM PINN FNO KM
11 $36.63$ $300.2$ $744.3$ $1.918$
$2^{2}$ $17.05$ $20.68$ $58.71$ $0.0028$
$4^{2}$ $7.478$ $8.654$ $37.68$ N/A
$N_{pde}=10^{2}$ $3.228$ $6.457$ $13.37$ N/A
average $16.10\pm 12.87$ $83.99\pm 124.9$ $213.5\pm 306.9$ $0.9603\pm 0.9576$
From Table LABEL:tab:advection-1d-error, we conclude that all solvers are sensitive to the number of training data, where larger $C_{\text{scale}}$ leads to better precision on test points. Kansa outperforms other methods in both accuracy and convergence speed, achieving the least error (up to $10^{-6}$) with only $C_{\text{scale}}=4^{2}$. However, due to the increasing computational cost above $C_{\text{scale}}=10^{2}$, memory limit was exceeded.
4.2.2 Inverse problem
For the 1D advection equation equation 45 initialized in Table LABEL:tab:advection-1d-setup, we set up the inverse PDE problem to infer the initial parameter $\beta$ from the observation data $u^{\text{obs}}$ at all time steps. All methods are evaluated at their best performance from the forward problem. The results are summarized in Table LABEL:tab:advection-1d-inverse-error, with the initial parameter $\beta_{0}$ set.
Table 4: Inverse predictions of $\beta$ on advection equation, where the ground truth $\beta=0.4$.
$\beta_{0}$ FDM PINN FNO KM $\beta_{0}$ FDM PINN FNO KM
$0.2$ $0.402$ $0.39987$ $0.3985$ $0.402$ $1.0$ $1.000$ $0.40446$ $1.2267$ $0.402$
For local optimization methods when searching for the optimal parameter, they stuck at different local minima depending on the initial guess $\beta_{0}$. With different runs of initial guesses, they give more precise predictions with more computational cost.
4.3 Extension 1: Kansa method for coupled PDEs
The Lotka-Volterra equations equation 47 are initialized as per Table LABEL:tab:lotka-volterra-setup, where the number of domain quadrature points $N_{\mathcal{R}}=100$, and initial condition points $N_{d}=1$. For evaluation, the test points $N_{\text{test}}=64$. The results from Kansa method are summarized in Table LABEL:tab:LV-maxwell-error, where the Gaussian RBF shape parameters, as defined in equation 3, are set as $\epsilon=0.2$ for both $x(t)$ and $y(t)$.
Table 5: 1D Lotka-Volterra equations experimental setup.
time range parameter initial conditions
$t_{0}=0,t_{f}=200$ $\alpha=0.1,\beta=0.02,\delta=0.01,\gamma=0.1$ $x(0)=40,y(0)=9$
The 1D Maxwell’s equations as defined in equation 58 are initialized per Table LABEL:tab:maxwell-1d-setup, where the speed of time propagation $c=1$, the number of domain quadrature points $N_{\mathcal{R}}=12\times 12$, and initial condition points $N_{d}=24$. For evaluation, the test points $N_{\text{test}}=10\times 10$. The shape parameter of Gaussian RBF, as defined in equation 3, is set as $\epsilon_{x}=0.21$ and $\epsilon_{y}=0.2$ for Lotka-Volterra equations and $\epsilon_{E}=16$ and $\epsilon_{B}=16$ for Maxwell’s equations, respectively.
Table 6: 1D Maxwell’s equations experimental setup.
domain time span parameter initial conditions
$x_{0}=0,$ $x_{f}=1$ $t_{0}=0,$ $t_{f}=\frac{1}{2}$ $c=1$ $E_{z}(x,0)=\sin(2\pi x)+\frac{1}{2}\sin(4\pi x)$ $B_{y}(x,0)=\cos(2\pi x)+\frac{1}{2}\cos(4\pi x)$
4.3.1 Forward problem
[…]
$C_{\text{scale}}$ $x(t)$ $y(t)$ $E_{z}(x,t)$ $B_{y}(x,t)$
11 $0.1279353$ $0.055667494$ 0.8049189 0.5894967
44 $0.04539858$ $0.06230465$ 0.4383743 0.3830594
Accuracy. The results from Kansa method are summarized in Table LABEL:tab:LV-maxwell-error. Both errors converge with increasing $C_{\text{scale}}$. Efficiency. The training time and inference time of Kansa method on Lotka-Volterra equations are $0.4034$ and $0.0001$ seconds, respectively. The training time and inference time of Kansa method on Maxwell’s equations are $0.4486$ and $0.0005$ seconds.
4.3.2 Inverse problem
For the Lotka-Volterra defined in equation 47 initialized in Table LABEL:tab:lotka-volterra-setup, we set up the inverse problem to infer the initial parameter $\alpha$, $\beta$, $\delta$ and $\gamma$ from observation $x^{\text{obs}}(t)$ and $y^{\text{obs}}(t)$ at all time steps.
Table 8: Inverse predictions of $\alpha$, $\beta$, $\delta$ and $\gamma$ on Lotka-Volterra equations.
$\alpha$ $\beta$ $\delta$ $\gamma$ $\alpha$ $\beta$ $\delta$ $\gamma$
reference $0.1$ $0.02$ $0.01$ $0.1$ prediction $0.102$ $0.0207$ $0.0100$ $0.0994$
With the initial guess all set to $1$, the results are summarized in Table LABEL:tab:lotka-volterra-inverse-acc. Despite the four-dimensional search space, the optimization algorithm SciPy Powell method successfully infers the parameters with high accuracy and decent computational cost.
4.4 Extension 2: Kansa method for nonlinear PDEs
The Burgers’ equation defined in equation 63 is initialized as per Table LABEL:tab:burgers-setup, where the number of domain quadrature points $N_{\mathcal{R}}=64\times 16$, i.e. initial condition (IC) points $N_{d}=64$ and the boundary condition (BC) points $N_{\mathcal{B}}=16\times 2$. For evaluation, the test points $N_{\text{test}}=48\times 12$. The Gaussian RBF shape parameter, as defined in equation 3, is set as $\epsilon=0.9$.
Table 9: Burgers’ equation experimental setup.
domain time span param. ICs BCs
$x_{0}=-10,x_{f}=10$ $t_{0}=0,t_{f}=4$ $\nu=0.5$ per equation 74 $u(x_{0})=1,u(x_{f})=0$
4.4.1 Forward problem
Stability. For four time discretization schemes, only forward Euler scheme is unstable (Table LABEL:tab:Stability-forward-euler), where time step $\Delta t$ exceeds the stability limit when $C^{t}_{\text{scale}}=1$ and $2$ according to CFL condition.
Table 10: Stability test of forward Euler Kansa method on Burgers’ equation.
$C^{t}_{\text{scale}}$ 1 2 4 10
[…]
Stability unstable unstable stable stable
Accuracy. From Table LABEL:tab:burgers-error, we observe that fully non-linear approach outperforms other time-stepping schemes. It’s hard to determine whether IMEX or backward Euler is more accurate theoretically. However, Crank-Nicolson scheme is definitely more accurate than both IMEX and backward Euler, since it’s second-order accurate in time while the other two are only first-order accurate.
[…]
forward IMEX backward Crank–Nicolson fully non-linear
$3.74\times 10^{31}$ $1.68$ $1.33$ $1.29$ $0.012$
Computational efficiency. We measure the training and inference time of different Kansa methods on Burgers’ equation in Table LABEL:tab:burgers-time. The non-linear solver used is the SciPy least-squares.
Table 12: Train or infer time of Burgers’ equation using Kansa methods (in seconds).
forward IMEX backward Crank–Nicolson fully non-linear
Training $0.34$ $0.47$ $2.33$ $1.66$ $99.2$
Inference $0.436$ $0.272$ $1.053$ $1.109$ $0.005$
Training time for fully nonlinear approach is longer because each step involves heavier computation with substantial memory (Table LABEL:tab:nonlinear-kansa-error-bound), further compounded by nonlinear solvers. Four time-stepping schemes have much less training time. The forward Euler is unstable when the stability condition is not satisfied. Inference time of fully non-linear approach is significantly reduced, due to the reuse of coefficient from the training phase. Despite a full test-time recomputation from scratch, the inference time of four time-stepping schemes remains acceptable for most practical applications.
4.4.2 Inverse problem
For the Burgers’ equation defined in equation 63 initialized in Table LABEL:tab:burgers-setup, we set up the inverse PDE problem to infer the initial parameter $\nu$ from the observation data $u^{\text{obs}}$ at all time steps. The results are summarized in Table LABEL:tab:burger-inverse-acc, with the initial parameter $\nu_{0}$ set as $0.1$.
Table 13: Inverse predictions of $\nu$ on Burgers’ equation, where the ground truth $\nu=0.5$.
forward IMEX backward Crank–Nicolson fully non-linear
$0.388$ $0.535$ $0.467$ $0.502$ $0.500$
Accuracy. Under same optimizer and initial guess, the Crank-Nicolson scheme confirms its theoretical advantage (Table LABEL:tab:nonlinear-kansa-error-bound) over both IMEX and backward Euler, which stuck at local minima. Computational efficiency. Fully non-linear approach requires retraining for each new parameter, which is computationally expensive (Table LABEL:tab:burgers-time). To speed up the per-run training time, it is trained with a maximum iteration. Stability. The forward Euler scheme is unstable when given large $\Delta t$.
5 Conclusions
This paper extends Zhong et al. (2023) RBF framework solver beyond original scope of linear PDEs. In particular, we generalize its PDE solver to handle multi-dependent-variable and nonlinear PDEs, addressing the loss property of linear reordering. These broaden the applicability of CNF-driven self-tuning mesh-free solvers to both forward modeling and inverse problem formulations.
In addition, this work contributes a systematic empirical study of how CNF solvers compare with established classical and neural PDE solvers. By implementing representative prior methods and evaluating them across benchmark problems, we assess their relative performance in terms of solution accuracy, efficiency, convergence and complexity. Such comparisons clarify the strengths and limitations of CNF-based approaches within the broader landscape of PDE solvers.
Overall, this paper demonstrates that learning-guided Kansa solvers can serve as a promising and flexible tool for coupled or nonlinear PDE systems. Future work includes theoretical analysis of error and convergence properties, application to neural field in computing, and integration with differentiable pipelines in scientific domains.
Appendix A Linear operator
A linear operator Iserles (2008) is a function $\mathcal{F}:V\rightarrow W$ that maps one vector space $V\in\mathbb{R}$ to another, or itself [^5] If the domain and codomain are the same vector space, i.e., $\mathcal{F}:V\rightarrow V$, it’s called a linear transformation or operator on $V$., $W\in\mathbb{R}$, and preserving the operations of vector addition and scalar multiplication, also known as homogeneity. Thus, for all vectors $\mathbf{u_{i}}\in V$ and all scalars $c$, the following features hold:
$\mathcal{F}(\sum_{i=1}^{n}\mathbf{u}_{i})=\sum_{i=1}^{n}\mathcal{F}(\mathbf{u}_{i}),\quad\text{vector additivity},$ (36)
$\mathcal{F}(c\cdot\mathbf{u})=c\cdot\mathcal{F}(\mathbf{u}),\quad\text{scalar multiplication}.$
Linear operators are fundamental in Linear Algebra for processing matrices, Quantum Mechanics for observables, Machine Learning, and Signal Processing. This forms the basis for Kansa method for linear PDEs.
Here are several commonly used examples of linear operators below, among which some are used in this work for PDE solver algorithms.
  • Matrix multiplication: For a matrix $A$, the function $A:\mathbb{R}^{n}\rightarrow\mathbb{R}^{m}$ is a linear operator,
$\mathcal{F}(\mathbf{x})=A\mathbf{x}.$ (37)
  • Integral operator: The operator that integrates a function over a fixed interval $[a,b]$ is a linear operator,
$I(f)=\int_{a}^{b}f(x)\,dx.$ (38)
  • Differentiation: The operator taking the derivative in a function space is a linear operator, because differentiation preserves addition and scalar multiplication,
$D_{x}(f)=\frac{\partial f}{\partial x}.$ (39)
  • Gradient operator: In multivariable calculus, the gradient operator $\nabla$ is a linear operator that maps a scalar field to a vector field,
$\nabla f=\left(\frac{\partial f}{\partial x},\frac{\partial f}{\partial y},\frac{\partial f}{\partial z}\right).$ (40)
  • Divergence operator: In vector calculus, the divergence operator $\nabla\cdot$ is a linear operator that maps a vector field to a scalar field,
$\nabla\cdot\mathbf{F}=\frac{\partial F_{x}}{\partial x}+\frac{\partial F_{y}}{\partial y}+\frac{\partial F_{z}}{\partial z}.$ (41)
  • Laplace operator: In the context of partial differential equations, the Laplace operator $\Delta$ is a linear operator that maps a scalar field to another scalar field,
$\Delta f=\nabla\cdot(\nabla f)=\nabla^{2}f=\frac{\partial^{2}f}{\partial x^{2}}+\frac{\partial^{2}f}{\partial y^{2}}+\frac{\partial^{2}f}{\partial z^{2}}.$ (42)
  • Curl operator: In vector calculus, the curl operator $\nabla\times$ is a linear operator that maps a vector field to another vector field,
$\nabla\times\mathbf{F}$ $=\left(\frac{\partial F_{z}}{\partial y}-\frac{\partial F_{y}}{\partial z},\frac{\partial F_{x}}{\partial z}-\frac{\partial F_{z}}{\partial x},\frac{\partial F_{y}}{\partial x}-\frac{\partial F_{x}}{\partial y}\right)$ $=\begin{vmatrix}\hat{i}&\hat{j}&\hat{k}\\ \frac{\partial}{\partial x}&\frac{\partial}{\partial y}&\frac{\partial}{\partial z}\\ F_{x}&F_{y}&F_{z}\end{vmatrix}$. (43)
Appendix B Partial differential equations
B.1 Boundary and initial conditions (BCs and ICs)
Since solution to differential equations contain integration constants, which is non-unique, additional conditions are required to enforce uniqueness. The boundary conditions (BCs) specify the function $u$ behavior on the domain boundary $\partial\Omega$, whereas the initial conditions (ICs) from time scale perspective are given at $t=0$. The formulation is defined in equation 1.
There are some common boundary conditions, defined over the boundary $\Omega=[x_{0},x_{f}]$ in 1D space, where $\{g_{i}\}_{i=1}^{4}$ are given closed-form functions,
Zero BC: $u(x_{0},t)=0,\;u(x_{f},t)=0,$ (44)
Dirichlet BC: $u(x_{0},t)=g_{1}(t),\;u(x_{f},t)=g_{2}(t),$
von Neumann BC: $\frac{\partial u}{\partial x}(x_{0},t)=g_{3}(t),\;\frac{\partial u}{\partial x}(x_{f},t)=g_{4}(t).$
B.2 Summary of PDEs
Table 14: Summary of PDEs with different characteristics.
Equation Domains Linearity Solution dim.
Advection Physics, Graphics Linear 1
Wave Physics, Graphics Linear 1
Lotka-Volterra Biology Linear 2
Maxwell Physics Linear 2
Burgers Physics, Graphics Nonlinear 1
B.3 1D Advection equation
The advection equation Takamoto et al. (2022) models the linear transport of a scalar quantity $u(x,t)$, which is changed over time $t$ and space $x$, as follows:
[…]
where parameter $\beta\in\mathbb{R}$ is the advection velocity, and $u_{0}(x)$ is the initial condition given at $t=0$. The analytical solution of equation 45 is,
$u(x,t)=u_{0}(x-\beta t).$ (46)
The positivity of parameter $\beta$ indicates the direction of wave propagation. From equation 46, when $\beta>0$, the wave propagates rightwards, and vice versa. The solution is visualized in Figure 1, with initial condition $u_{0}(x)=\sin(2\pi x)$, $x\in[0,1]$.
Figure 1: Advection equation solution visualization in 1D, 2D and 3D.
B.4 Lotka-Volterra predator-prey model
Lotka-Volterra predator-prey model Bacaër (2011) relates the populations of prey $x(t)$ and predators $y(t)$ at time $t$ in a dynamic biological system via coupled differential equations, also applicable to other fields, e.g. the unemployment rate with respect to wage growth Orlando and Sportelli (2021) and many more,
$\begin{cases}x^{\prime}(t):=\frac{dx(t)}{dt}=\alpha x(t)-\beta x(t)\cdot y(t),\\ y^{\prime}(t):=\frac{dy(t)}{dt}=\delta x(t)\cdot y(t)-\gamma y(t).\end{cases},t\in[t_{0},t_{f}].$ (47)
where $\alpha$ is the prey growth rate, $\beta$ is the predation rate, $\delta$ is the ratio of neonate predators to eaten prey, and $\gamma$ is the predator death rate. It assumes that there would be unlimited food supply for the prey, and thus exponential growth $\alpha x(t)$. The multiplicative term $x(t)\cdot y(t)$ represents the encounters between prey and predators statistically.
The system has no explicit analytical solution, but the implicit solution exists. After scaling of variables,
$x^{\ast}(t)=\frac{\delta}{\gamma}x(t),\quad y^{\ast}(t)=\frac{\beta}{\alpha}y(t),\quad\tau=\alpha t,$ (48)
By plugging into equation 47, and dividing the first equation by the second,
$\frac{dy^{\ast}}{dx^{\ast}}=\frac{\gamma}{\alpha}\cdot\frac{y^{\ast}(x^{\ast}-1)}{x^{\ast}(y^{\ast}-1)},$ (49)
The implicit solution is given by integration separation of variables, for which $C_{\text{L-V}}\in\mathbb{R}$ is the integration constant,
$\ln(y^{\ast})-y^{\ast}-\frac{\gamma}{\alpha}[\ln(x^{\ast})-x^{\ast}]=C_{\text{L-V}}.$ (50)
Figure 2 shows the solution with phase space given by the above implicit solution, which depends on the initial conditions $x(0)$ and $y(0)$.
(a) Prey and predator solution populations
(b) Phase space trajectory
Figure 2: Lotka-Volterra predator-prey model solution and phase space.
B.5 Maxwell’s equations
In electromagnetism, Maxwell’s equations Greiner (1998) relate the electric field $\mathbf{E}(\mathbf{r},t)$ and magnetic field $\mathbf{B}(\mathbf{r},t)$ with spatial position $\mathbf{r}\in\mathbb{R}^{3}$ and time $t$, to the electric charge density $\rho(\mathbf{r},t)\in\mathbb{R}$ and current density $\mathbf{J}(\mathbf{r},t)\in\mathbb{R}^{3}$. The differential form is as follows,
[…]
where constants $\mu_{0},\epsilon_{0}\in\mathbb{R}^{+}$ are the vacuum permeability and permittivity respectively. Their product is the reciprocal of the square of the speed of light $c\approx 3\times 10^{8}$ $\mathrm{m\,s^{-1}}$ in vacuum,
$\mu_{0}\epsilon_{0}=\frac{1}{c^{2}}.$ (52)
The first two equations state that the electric field $\mathbf{E}$ sourced by electric charges, and no magnetic monopoles exist. The last two equations depict how a time-varying magnetic field $\mathbf{B}$ induces an electric field $\mathbf{E}$, and vice versa with the addition of current density $\mathbf{J}$. In general, Maxwell’s equations are linear with respect to $\mathbf{E}$ and $\mathbf{B}$.
Taking the curl of Faraday’s law and Ampère-Maxwell law respectively,
$\begin{cases}\nabla\times(\nabla\times\mathbf{E})=-\frac{\partial}{\partial t}(\nabla\times\mathbf{B}),\\ \nabla\times(\nabla\times\mathbf{B})=\mu_{0}(\nabla\times\mathbf{J})+\mu_{0}\epsilon_{0}\frac{\partial}{\partial t}(\nabla\times\mathbf{E}).\end{cases}$ (53)
By vector calculus identity of $\nabla\times(\nabla\times\mathbf{E})=\nabla(\nabla\cdot\mathbf{E})-\Delta\mathbf{E}$, and plugging in Ampère-Maxwell law on the right-hand side of the first equation,
$\begin{cases}\nabla(\nabla\cdot\mathbf{E})-\Delta\mathbf{E}=-\mu_{0}\frac{\partial\mathbf{J}}{\partial t}-\mu_{0}\epsilon_{0}\frac{\partial^{2}\mathbf{E}}{\partial t^{2}},\\ \nabla(\nabla\cdot\mathbf{B})-\Delta\mathbf{B}=\mu_{0}(\nabla\times\mathbf{J})+\mu_{0}\epsilon_{0}\frac{\partial}{\partial t}(\nabla\times\mathbf{E}).\end{cases}$ (54)
By substituting Gauss’s law for $\nabla\cdot\mathbf{E}$ in the first equation, Gauss’s law for magnetism for $\nabla\cdot\mathbf{B}$ and Faraday’s law for $\nabla\times\mathbf{E}$ in the second equation, the two equations after rearrangement are inhomogeneous, i.e. including source terms $\mathbf{F}(\mathbf{r},t)$, wave equations, taking the forms of $c^{2}\Delta u-u_{tt}=\mathbf{F}$,
$\begin{cases}\Delta\mathbf{E}-\mu_{0}\epsilon_{0}\frac{\partial^{2}\mathbf{E}}{\partial t^{2}}=\nabla(\frac{\rho}{\epsilon_{0}})+\mu_{0}\frac{\partial\mathbf{J}}{\partial t},\\ \Delta\mathbf{B}-\mu_{0}\epsilon_{0}\frac{\partial^{2}\mathbf{B}}{\partial t^{2}}=-\mu_{0}(\nabla\times\mathbf{J}).\end{cases}$ (55)
To simplify the problem, we take the one-dimensional (1D) electromagnetic wave propagating along $x$-axis without sources, i.e. $\rho=0$ and $\mathbf{J}=0$, with the electric field $\mathbf{E}(\mathbf{r},t)=(0,0,E_{z}(x,t))$ along $z$-axis and magnetic field $\mathbf{B}(\mathbf{r},t)=(0,B_{y}(x,t),0)$ along $y$-axis respectively. By expanding the defintion of curl $\nabla\times$ operators, and removing the zero terms,
$\nabla\times\mathbf{E}=\left(\frac{\partial E_{z}}{\partial y}-\frac{\partial E_{y}}{\partial z},\frac{\partial E_{x}}{\partial z}-\frac{\partial E_{z}}{\partial x},\frac{\partial E_{y}}{\partial x}-\frac{\partial E_{x}}{\partial y}\right)=\left(0,-\frac{\partial E_{z}}{\partial x},0\right),$ (56)
thus the reduced last two equations of Maxwell’s equations equation 51 are,
$\begin{cases}\frac{\partial E_{z}}{\partial x}=-\frac{\partial B_{y}}{\partial t},\\ \frac{\partial B_{y}}{\partial x}=-\mu_{0}\epsilon_{0}\frac{\partial E_{z}}{\partial t}.\end{cases}$ (57)
By taking partial derivatives $\partial_{x}$ and $\partial_{t}$, and simplifying, the 1D wave solutions are [^6] The 1D wave equation aligned with the general inhomogeneous wave equation equation 55, with $\mathbf{F}=0$.,
$\begin{cases}\frac{\partial^{2}E_{z}}{\partial x^{2}}-\mu_{0}\epsilon_{0}\frac{\partial^{2}E_{z}}{\partial t^{2}}=0,\\ \frac{\partial^{2}B_{y}}{\partial x^{2}}-\mu_{0}\epsilon_{0}\frac{\partial^{2}B_{y}}{\partial t^{2}}=0.\end{cases}$ (58)
The initial conditions at $t=0$ are given as follows,
$E_{z}(x,0)=f(x),\quad B_{y}(x,0)=g(x),\quad x\in[x_{0},x_{f}].$ (59)
Let $u=E_{z}+B_{y}$ and $v=E_{z}-B_{y}$ and with change of variables $x_{t}=x_{t=0}\pm ct$, where $c$ defined in equation 52 is the speed of light in vacuum. According to d’Alembert’s formula le Rond D’Alembert (1747),
$\begin{cases}u(x,t)=u(x-ct,0)=f(x-ct)+g(x-ct),\\ v(x,t)=v(x+ct,0)=f(x+ct)-g(x+ct).\end{cases}$ (60)
By reversing the change of variables, the analytical solutions to equation 58 are,
$\begin{cases}E_{z}=\frac{1}{2}(u+v)=\frac{1}{2}[f(x-ct)+f(x+ct)]+\frac{1}{2}[g(x-ct)-g(x+ct)],\\ B_{y}=\frac{1}{2}(u-v)=\frac{1}{2}[f(x-ct)-f(x+ct)]+\frac{1}{2}[g(x-ct)+g(x+ct)].\end{cases}$ (61)
The solution is visualized in Figure 3, with initial condition given as,
$f(x)=\sin(2\pi x)+0.5\sin(4\pi x),\;g(x)=\cos(2\pi x)+0.5\cos(4\pi x)\quad x\in[0,1],t\in[0,0.5].$ (62)
Figure 3: Maxwell’s equations solution visualization in 1D, 2D and 3D.
B.6 Viscous Burgers’ equation
Viscous Burgers’ equation Takamoto et al. (2022) captures both non-linear advection, also known as convection and diffusion phenomena in dynamics,
[…]
where viscosity $\nu\in\mathbb{R}^{+}$ is the positive constant, and $u_{0}(x)$ is the initial condition given at $t=0$. By Cole-Hopf transformation Hopf (1950), unknown function $u(x,t)$ is converted into $\phi(x,t)$ via,
$u(x,t)=-2\nu\frac{\partial}{\partial x}\ln\phi(x,t)=-2\nu\frac{1}{\phi(x,t)}\frac{\partial\phi(x,t)}{\partial x}\equiv-2\nu\frac{\phi_{x}}{\phi}.$ (64)
By chain rule and quotient rule of differentiation, the first-order and second-order spatial or temporal derivatives of $u(x,t)$ are,
$\frac{\partial u(x,t)}{\partial x}=2\nu\left(\frac{\phi_{x}^{2}}{\phi^{2}}-\frac{\phi_{xx}}{\phi}\right),\quad\frac{\partial u(x,t)}{\partial t}=2\nu\left(\frac{\phi_{x}\phi_{t}}{\phi^{2}}-\frac{\phi_{xt}}{\phi}\right),\quad$ (65)
$\frac{\partial^{2}u(x,t)}{\partial x^{2}}=2\nu\left(\frac{3\phi_{x}\phi_{xx}}{\phi^{2}}-\frac{2\phi_{x}^{3}}{\phi^{3}}-\frac{\phi_{xxx}}{\phi}\right).$
By plugging equation 65 into equation 63 and simplifying,
$2\nu\left(\frac{\phi_{x}\phi_{t}}{\phi^{2}}-\frac{\phi_{xt}}{\phi}-\nu\frac{\phi_{x}\phi_{xx}}{\phi^{2}}+\nu\frac{\phi_{xxx}}{\phi}\right)=0,\quad x\in[x_{0},x_{f}],t\in[t_{0},t_{f}],$ (66)
With the inversion of quotient rule, equation 66 is rearranged as,
$2\nu\frac{\partial}{\partial x}\left(\frac{\nu\phi_{xx}-\phi_{t}}{\phi}\right)=0,\quad x\in[x_{0},x_{f}],t\in[t_{0},t_{f}].$ (67)
By integrating equation 67 with respect to $x$ and introducing an integration function $f(t)$,
$\frac{\nu\phi_{xx}-\phi_{t}}{\phi}=f(t),\quad x\in[x_{0},x_{f}],t\in[t_{0},t_{f}].$ (68)
Now introduce $f(t)=\frac{dF(t)}{dt}$ and $\tilde{\phi}=\phi\cdot e^{F(t)}$, thus the derivatives of $\tilde{\phi}$ are,
$\frac{\partial\tilde{\phi}}{\partial t}=e^{F(t)}\left(\phi_{t}+\phi\frac{dF(t)}{dt}\right),\quad\frac{\partial^{2}\tilde{\phi}}{\partial x^{2}}=e^{F(t)}\phi_{xx},$ (69)
by plugging them into equation 68. The resulting equation is reduced to the standard heat equation,
$\nu\frac{\partial^{2}\tilde{\phi}(x,t)}{\partial x^{2}}-\frac{\partial\tilde{\phi}(x,t)}{\partial t}=0,\quad x\in[x_{0},x_{f}],t\in[t_{0},t_{f}].$ (70)
The solution of equation 70 is formed by heat kernel $\Phi(x,t)$ convolved with the initial condition $\tilde{\phi}_{0}(x)=\tilde{\phi}(x,0)$ Evans (2010),
$\tilde{\phi}(x,t)=\int_{-\infty}^{\infty}\Phi(x-x^{\prime},t)\tilde{\phi}_{0}(x^{\prime})\;dx^{\prime},\quad\text{where}\quad\Phi(x,t)=\frac{1}{\sqrt{4\pi\nu t}}e^{-\frac{x^{2}}{4\nu t}}.$ (71)
Note that the transformation from $\phi$ to $\tilde{\phi}$ does not change the Cole-Hopf transformation equation 64, since the additional multiplicative term $e^{F(t)}$ is independent of $x$,
$u(x,t)=-2\nu\frac{\partial}{\partial x}\ln\phi(x,t)=-2\nu\frac{\partial}{\partial x}\ln\tilde{\phi}(x,t).$ (72)
From the Cole-Hopf equation 72 at $t=0$ and via integration, the initial condition for $\tilde{\phi}(x,0)$ is thus,
$\tilde{\phi}_{0}(x)=-\frac{1}{2\nu}\int_{0}^{x}u_{0}(x^{\prime})\;dx^{\prime}.$ (73)
The analytical solution of equation 63 is thus plugging equation 73 into equation 71 and then into equation 72.
In this work, we consider when $u_{0}(-\infty)$ and $u_{0}(\infty)$ exist and $u_{0}^{\prime}(x)<0$ for all $x\in\mathbb{R}$, the explicit expression BATEMAN (1915) is then a steadily propagating wave as below,
$u(x,t)=c-\Delta u_{0}\tanh\left(\frac{\Delta u_{0}}{2\nu}(x-ct)\right),\quad\text{where }c=\frac{u_{0}(-\infty)+u_{0}(\infty)}{2},\Delta u_{0}=\frac{u_{0}(-\infty)-u_{0}(\infty)}{2}.$ (74)
The solution is visualized in Figure 4, with initial condition set by equation 74, $u_{0}(-\infty)=1$, $u_{0}(\infty)=0$, and $\nu=0.5$.
Figure 4: Burgers’ equation solution visualization in 1D, 2D and 3D.
B.7 PDE solvers
There are many attempts to solve the PDE solution field $u$. Among which, we categorize PDE solvers into two types, i.e. numerical analysis methods and neural-based methods.
Constrained optimization. In PDE solvers, constraints are boundary conditions, initial conditions, or PDE residuals. They can be in the form of either soft or hard constraints. The former is the cost functions that are penalised, while the latter is that can not be violated, e.g. (in)equality forms.
Table 15: Summary of different PDE solvers.
method motivation training supervised constraint
FDM grid-based $\times$ N/A hard
PINN physics-driven $\checkmark$ $\times$ soft
NO data-driven $\checkmark$ $\checkmark$ soft
- FNO data-driven $\checkmark$ $\checkmark$ soft
- PINO hybrid $\checkmark$ $\checkmark$ soft
KM mesh-free grid $\times$ N/A hard
- CNF KM on neural fields $\checkmark$ $\times$ hard
B.7.1 Finite difference method (FDM)
Considering a discretized sequence $\mathbf{u}\in\mathbb{R}^{N\times M}$ of the continuous function $u(x,t)$ as in equation 45. Along the spatial dimension $x$ and temporal dimension $t$, there are $N$ and $M$ sampled points respectively. The finite difference operators Iserles (2008) defined on per element, are as follows:
$(\Delta\mathbf{u})_{i}=\begin{cases}\begin{aligned} (\Delta^{+}\mathbf{u})_{i}&=u_{i+1}^{j}-u_{i}^{j},&\quad\text{forward difference.}\\ (\Delta^{-}\mathbf{u})_{i}&=u_{i}^{j}-u_{i-1}^{j},&\quad\text{backward difference.}\\ (\Delta^{0}\mathbf{u})_{i}&=\frac{u_{i+1}^{j}-u_{i-1}^{j}}{2},&\quad\text{central difference.}\end{aligned}\end{cases},$ (75)
for which $i\in\{0,1,\ldots,N-1\}$ is the spatial index, and $j\in\{0,1,\ldots,M-1\}$ is the temporal index of the sequence $\mathbf{u}$.
The partial equations often involve full and/or partial derivatives, where differential operators can be discretized into difference operators equation 75 via finite difference method Bargteil and Shinar (2018). By Taylor expansion of $u(x\pm\Delta x,t)$ around $u(x,t)$ up to the first order error, the corresponding examples for the spatial derivative are,
[…]
$=\frac{(\Delta^{0}\mathbf{u})_{i}}{\Delta x}+O(\Delta x^{2})\approx\frac{(\Delta^{0}\mathbf{u})_{i}}{\Delta x}=\frac{u_{i+1}^{j}-u_{i-1}^{j}}{2\Delta x},$ central difference.
where $\Delta x$ is the spatial spacing, with spatial index $i$ and temporal index $j$ defined above. The upwind scheme Patankar (1980) considers where the information comes from, e.g. when $\beta>0$, the wave propagates rightwards, and thus $u_{i}^{j}$ is influenced by $u_{i-1}^{j}$, and vice versa for downwind scheme. Under the upwind scheme, the advection equation equation 45 is therefore as the following ODE,
$\frac{\partial u(x,t)}{\partial t}+\beta[\mathbb{I}_{\beta>0}\frac{(\Delta^{-}\mathbf{u})_{i}}{\Delta x}+\mathbb{I}_{\beta<0}\frac{(\Delta^{+}\mathbf{u})_{i}}{\Delta x}]=0,$ (77)
where the indicator function $\mathbb{I}_{\beta>0}=\begin{cases}1&\text{if }\beta>0,\\ 0&\text{otherwise.}\end{cases}$ is for controlling different cases of $\beta$ [^7] Alternatively, one may use $\max(\beta,0)$ and $\min(\beta,0)$ to replace $\beta\mathbb{I}_{\beta>0}$ and $\beta\mathbb{I}_{\beta<0}$ respectively..
By forward Euler method for ODEs, the temporal derivative is discretized via the forward difference operator. After which, the advection equation equation 45 is simplified as, with $\Delta t$ being the temporal spacing,
$\frac{u_{i}^{j+1}-u_{i}^{j}}{\Delta t}+\beta[\mathbb{I}_{\beta>0}\frac{(\Delta^{-}\mathbf{u})_{i}}{\Delta x}+\mathbb{I}_{\beta<0}\frac{(\Delta^{+}\mathbf{u})_{i}}{\Delta x}]=0.$ (78)
With algebraic reordering, the upwind scheme update rule is thus,
$u_{i}^{j+1}=u_{i}^{j}-\frac{\beta\Delta t}{\Delta x}[\mathbb{I}_{\beta>0}(\Delta^{-}\mathbf{u})_{i}+\mathbb{I}_{\beta<0}(\Delta^{+}\mathbf{u})_{i}].$ (79)
Stability condition. For implicit numerical schemes, e.g. the backward Euler method, the solution is unconditionally stable. However, for explicit numerical schemes, e.g. the forward Euler method above, stability conditions must be satisfied to avoid numerical instability, which we briefly introduce below.
In the 1D space, the scalar Courant number $C$, also known as the CFL stability criteria, measures the ratio of how far the wave propagates in one time interval $\Delta t$ to the spatial spacing $\Delta x$. The CFL condition Courant et al. (1928) states that $C$ must satisfy,
$C=\frac{|\beta|\Delta t}{\Delta x}\leq C_{max},$ (80)
where $C_{max}$ is a problem-dependent constant. It sets the maximum allowable time step $\Delta t$ for a given $\Delta x$, for numerical stability.
B.7.2 Physics-informed neural network (PINN)
Physics-informed neural network (PINN) Raissi et al. (2019) is a data-driven approach for functional PDE approximation, which requires a large labeled dataset but has the ability to generalize. Consider the general form of PDEs defined in equation 1, PINNs approximate the unknown solution $u(x,t)\in\mathcal{U}$ with a neural network $\hat{u}_{\theta}(x,t)\in\mathcal{U}$, i.e. $\hat{u}_{\theta}(x,t)\approx u(x,t)$, parameterized by updatable parameters $\theta\in\Theta$.
Residual $\mathcal{R}_{\theta}$ of the PDEs is calculated without supervised data for the neural network $\hat{u}_{\theta}$, which is minimized via automatic differentiation Baydin et al. (2018) [^8] Example of $\frac{\partial u}{\partial x}$: u_x = torch.autograd.grad(outputs=u, inputs=x, create_graph=True)[0] during training for generalizability,
$\mathcal{R}_{\theta}(x,t)\in\mathcal{Y}=\mathcal{D}[\hat{u}_{\theta}](x,t)-f(x,t),\quad x\in\Omega,t\in[t_{0},t_{f}].$ (81)
The residual loss $L_{R}$ [^9] Note that $L_{R}$ is the same as the risk $\mathcal{R}$ defined in equation 97, but for the residual $\mathcal{R}_{\theta}$ instead of the solution $u$., also known as the physics-informed loss, is defined to be the $p$-norm of the residual $\mathcal{R}_{\theta}$ in equation 81. During training, $N_{\mathcal{R}}$ quadrature points are sampled, where the integral loss is approximated by the discretized loss $\mathcal{L}_{R}$ with weights $\omega_{k}$ at each sample index $k$ and training error $\mathcal{E}_{T}(\theta)$,
[…]
$\text{By quadrature,}\quad=$ $\underbrace{\sum_{k=1}^{N_{\mathcal{R}}}\omega_{k}|\mathcal{R}_{\theta}(x_{k},t_{k})|^{p}}_{\text{discretized }\mathcal{L}_{R}}+\mathcal{E}_{T}(\theta)\approx\mathcal{L}_{R},\quad\text{where }\mathcal{E}_{T}(\theta)=L_{R}-\mathcal{L}_{R}.$
If considering the boundary conditions, the residual for the $i$-th boundary condition $\mathcal{R}^{\mathcal{B}_{i}}_{\theta}$ is calculated via equation 1 as well, after which the boundary condition loss $\mathcal{L}_{\text{BC}}$ is defined accordingly,
$\mathcal{R}^{\mathcal{B}_{i}}_{\theta}(x,t)\in\mathcal{Z}_{i}=\mathcal{B}_{i}[\hat{u}_{\theta}](x,t)-g_{i}(x,t),\quad x\in\partial\Omega_{i},t\in[t_{0},t_{f}].$ (83)
As defined in equation 99, the total error between the optimal solution from the network $\hat{u}_{\theta}$ and the ground truth $u$ is, by expanding equation 97,
[…]
During training, the network is optimized on supervised dataset $\{(x_{n},t_{n}),u(x_{n},t_{n})\}_{n=1}^{N_{d}}$, with $N_{d}$ being the total number of data. The supervised loss $\mathcal{L}_{\text{data}}$ [^10] Note that when the supervised data is only sampled on the boundary, the supervised loss and the boundary condition loss are the same. approximates the total error equation 84,
$\mathcal{L}_{\text{data}}=\frac{1}{N_{d}}\sum_{n=1}^{N_{d}}(|\hat{u}_{\theta}(x_{n},t_{n})-u(x_{n},t_{n})|^{p}).$ (85)
Training. PINN approximates the solution as $\hat{u}_{\theta}=u_{\theta^{\mathrm{opt}}}(x,t)$. To avoid overfitting due to the limited supervised data, the main goal is to minimize the unsupervised residual error $\mathcal{L_{R}}$ equation 82. With the addition of the supervised loss equation 85 and the boundary condition residual equation 83, the optimized theta is $\theta^{\mathrm{opt}}\approx\arg\min_{\theta\in\Theta}\mathcal{L}$, where the total training loss $\mathcal{L}_{\text{PINN}}$ is,
$\mathcal{L}_{\text{PINN}}=\underbrace{\sum_{k=1}^{N_{\mathcal{R}}}\omega_{k}|\mathcal{R}_{\theta}(x_{k},t_{k})|^{p}}_{\text{Discretized residual loss }\mathcal{L_{R}}}+\lambda_{1}\underbrace{\frac{1}{N_{d}}\sum_{n=1}^{N_{d}}(|\hat{u}_{\theta}(x_{n},t_{n})-u(x_{n},t_{n})|^{p})}_{\text{Supervised loss }\mathcal{L}_{\text{data}}}+\lambda_{2}\underbrace{\sum_{i}\sum_{b=1}^{N_{\mathcal{B}_{i}}}\omega_{b}^{\mathcal{B}_{i}}|\mathcal{R}^{\mathcal{B}_{i}}_{\theta}(x_{b},t_{b})|^{p}}_{\text{BC loss }\mathcal{L}_{\text{BC}}},$ (86)
with weights $\omega_{b}^{\mathcal{B}_{i}}$ at each sample index $b$ for $i$-th boundary condition and regularization parameters $\lambda_{1},\lambda_{2}>0$ for combining different losses.
Algorithm 1 Physics-Informed Neural Network training pseudocode.
1:Input: Initial parameters $\theta$ for network $\hat{u}_{\theta}$.
2:Output: Optimized parameters $\theta^{\mathrm{opt}}$ for network $\hat{u}_{\theta}$.
3:Hyperparameters: Learning rate $\eta$, number of training iterations $N_{\text{iter}}$.
4:while number of iterations $N_{\text{iter}}$ do
5:  Sample PDE points $x_{k}\in\Omega,t_{k}\in[t_{0},t_{f}]$ and boundary points $x_{b}\in\partial\Omega_{i},t_{b}\in[t_{0},t_{f}]$.
[…]
7:  Compute loss $\mathcal{L}_{\text{PINN}}=\mathcal{L_{R}}+\mathcal{L}_{\text{data}}+\mathcal{L}_{\text{BC}}$ by equation 86.
8:  By gradient descent, update $\theta\leftarrow\theta-\eta\,\nabla_{\theta}\mathcal{L}$.
9:end while
B.7.3 Neural operator (NO)
Operator learning. From the general form of PDEs equation 1, we assume that within the $\mathcal{D}$ operator, the source function $f$ or the initial conditions, there is a parameter $a$ of the same dimension as the solution $u$. In this subsection, we denote the differential operator $\mathcal{D}$ as $\mathcal{D}_{a}$, where the PDE is thus $\mathcal{D}_{a}[u]=f$.
Given the dataset $\{(a_{i}^{j},f_{i}^{j}),u_{i}^{j}|i=1,...,N_{\mathcal{R}}\}_{j=1}^{N_{pde}}$ with $N_{pde}$ PDE instances each $N_{\mathcal{R}}$ quadrature points, the idea of operator learning Li et al. (2020) is to learn the operator $\mathcal{G}$ mapping input $a\in\mathcal{A}:\mathbb{D}\to\mathbb{R}$ to the solution $u\in\mathcal{U}:\mathbb{D}\to\mathbb{R}$, i.e. $\mathcal{G}(a,f)=u$ connecting two function spaces $\mathcal{A}$ and $\mathcal{U}$ with infinite dimensions, which is challenging for neural networks since they are for finite dimensions instead.
Solution A. To solve this challenge, one solution is to parameterize the PDE $u=u(t,x,\mu)$, assuming that $a$ is measureable in finite dimension, i.e. $a=a(\mu),\mu\in\mathbb{R}^{d_{y}}$. It’s a techinique widely used in aircraft design and manufacturing Athanasopoulos et al. (2009), image processing Pérez et al. (2003). The training process is therefore to minimize the supervised loss $\mathcal{L}_{data}$ from data with $p$-norm, which measures how much the predicted solution $\mathcal{G}_{\theta}(a_{i},f_{i})$ deviates from the ground truth $u_{i}$ for each data point $i$,
[…]
The approximated solution is therefore $\hat{u}_{\theta}=\mathcal{G}_{\theta^{\mathrm{opt}}}(a,f)$, where $\theta^{\mathrm{opt}}\approx\arg\min_{\theta\in\Theta}\mathcal{L}_{data}$. There is no addition of the PDE residual loss $\mathcal{L_{R}}$ equation 81 as in PINNs for basic operator learning. An extension, termed as physics-informed neural operator (PINO) Li et al. (2024), combines the supervised loss $\mathcal{L}_{data}$ and the residual loss $\mathcal{L_{R}}$ as the total training loss,
[…]
Despite the simplicity in ideas, the parameterization suffers from how to sample from the given space, non-uniqueness, and low generalization to unseen $a$.
Solution B. Interpolation from the discretized grid, including a neural network based interpolator or traditional methods (linear, cubic, spline, etc). However, it suffers from inconsistency between the discretized and continuous functions.
Solution C. Generalize the neural network from discrete to continuous function space.
$(\mathcal{N}_{l}v)(x)=\sigma[A_{l}v(x)+B_{l}(x)+\int_{D}K_{l}(x,y)v(y)\;dy],\quad x\in D.$ (89)
Fast implemenentation via FFT.
Appendix C Architecture details
We adopt a multi-scale feed-forward neural network architecture for PINN Wang et al. (2021), which augments the original feed-forward architecture with input encoding layers of multiple frequency scales. There are three hidden layers each with 64 neurons. For FNO, we adopt the architecture with 16 modes retained in the spectral convolution layer, and the latent feature dimension is 64.
Appendix D Environment setup
All the measurements are conducted on a Mac M1, with a single-core CPU running at 3.2 GHz. In the following sections, the data points are uniformly sampled unless otherwise specified.
Appendix E Learning theory
E.1 Functional analysis
We introduce basic functional analysis concepts here for PDE solvers and later learning theory (Appendix § E).
The $p$-norm of a function $f:\Omega\subseteq\mathbb{R}^{d}\to\mathbb{R}$ is defined as,
$f||_{p}=(\int_{\Omega}|f(x)|^{p}\;dx)^{\frac{1}{p}},\text{for }1\leq p<\infty.$ (90)
The $p$-integrable function $f\in L^{p}(\Omega)$, is defined as
($f||_{p})^{p}=\int_{\Omega}|f(x)|^{p}\;dx<\infty. ($91)
For additional concepts used for learning theory, please refer to Appendix § E.2.
E.2 Functional analysis addendum
Smoothness Rudin (1976) of a function is defined to be the number of continuous derivatives it has. The class of function $f$ with smoothness $k\in\mathbb{N}^{+}$ has at least a $k$-th derivative, and is denoted as $f\in C^{k}$,
$C^{k}(\Omega)=\{f:\Omega\subseteq\mathbb{R}^{d}\to\mathbb{R}\;|\;\forall\alpha\leq k.\;\partial^{\alpha}f\text{ exists and is continuous}\}.$ (92)
When $k=\infty$, the function is differentiable at all orders. While not every function is not smooth, there is a generalization of smooth functions, i.e. Sobolev functions.
The weak derivative $f^{\prime}$ generalizes to include functions that are not differentiable, but locally integrable on bounded domain $[a,b]$. The $f^{\prime}$ definition is for all smooth test functions $\phi$, with $\phi(a)=\phi(b)=0$,
$\int_{a}^{b}f(x)\phi^{\prime}(x)\;dx$ $=[f(x)\phi(x)]_{a}^{b}-\int_{a}^{b}f^{\prime}(x)\phi(x)\;dx,\quad\text{by integration by parts},$ (93)
$=-\int_{a}^{b}f^{\prime}(x)\phi(x)\;dx,\quad\text{as }\phi(a)=\phi(b)=0.$
Sobolev spaces Adams and Fournier (2003) $W^{k,p}(\Omega)$ is a function space where all functions $f$ having weak derivatives up to order $k$ and every derivate is $p$-integrable via equation 91,
$W^{k,p}(\Omega)\subset L^{p}(\Omega)=\{f:\Omega\subseteq\mathbb{R}^{d}\to\mathbb{R}\;|\;\forall\alpha\leq k.\;\exists\;\partial^{\alpha}f\in L^{p}(\Omega)\},$ (94)
When $k=2$, it forms a Hilbert space, i.e. $W^{k,2}(\Omega)=H^{k}(\Omega)$.
E.3 Approximation theory of neural networks
We quote some known bounds for neural networks from theoretical machine learning field here, which are relevant to PDE solvers analysis later.
Universal approximation theorem Hornik et al. (1989) states that neural networks $\hat{u}_{\theta}$, for which parameters $\theta\in\Theta$, can approximate any continuous functions $u:\mathbb{R}^{d}\to\mathbb{R}$ with little error $\epsilon>0$ in the $p$-norm of function space $\mathcal{U}$, with an extension to their differential operator $\mathcal{D}$,
[…]
Optimal DNN functions approximation theorem Yarotsky (2018). Assuming a continuous function $u\in W^{s,p}$ as defined in equation 94, where $s\in\mathbb{N}^{+}$ is the smoothness of $u$, there exists a neural network $\hat{u}_{\theta}$ with $M$ parameters, such that the error bound is,
$\hat{u}_{\theta}-u||_{p}=O(M^{-\frac{s}{d}}),$ (96)
where $d$ is the input dimension of $u$. It means intuitively that the smoother and lower-dimensional the $u$ is, the easier for it to be approximated by a neural network $\hat{u}_{\theta}$. For a fixed error $\epsilon$, the required number of parameters is $M=O(\epsilon^{-\frac{d}{s}})$, which suffers from the exponential growth of $d$, i.e. the curse of dimensionality Bellman (1957).
E.4 Error analysis
Error and risk estimation. Define the risk of the approximation $\hat{u}_{\theta}$ against the ground truth function $u:\Omega\to\mathbb{R}$ with $p$-norm integral,
[…]
For discretized computation, quadrature $\hat{\mathcal{R}}$ is used to approximate the integral risk $\mathcal{R}$ with $N$ sample points from the dataset, where $\omega_{k}$ is the weight at each sample index $k$,
[…]
where the training, also known as generalization or out-of-sample, error $\mathcal{E}_{T}(\theta)$ measures the difference between the integral risk and the discretized risk due to quadrature.
[…]
$\mathcal{E}(\theta):=\mathcal{R}(\hat{u}_{\theta})\leq\underbrace{\inf_{\theta^{\star}\in\Theta}\mathcal{R}(u_{\theta^{\star}})}_{\mathcal{E}_{A}(\theta)}+\underbrace{\hat{\mathcal{R}}(\hat{u}_{\theta})-\inf_{\theta^{\star}\in\Theta}\mathcal{R}(u_{\theta^{\star}})}_{\mathcal{E}_{O}(\theta)}+\underbrace{\mathcal{R}(\hat{u}_{\theta})-\hat{\mathcal{R}}(\hat{u}_{\theta})}_{\mathcal{E}_{T}(\theta)},$ (99)
the approximation error $\mathcal{E}_{A}$ measures the risk between the best network approximation $u_{\theta^{\star}}$ and ground truth $u$, optimization error $\mathcal{E}_{O}$ measures the trained network result $\hat{u}_{\theta}$ deviation from the best network approximation, and training error $\mathcal{E}_{T}$ defined in equation 98.
E.5 PINN learning theory
We briefly analyze the PINN error bound [^12] The PDE residual is considered here, whereas boundary and initial conditions are omitted for simplicity.. The total error between the optimal solution $\hat{u}_{\theta}$ and the ground truth $u$ is shown in equation 84. However, during training, the network doesn’t have access to the exact ground truth for $u$. Therefore, we aim to reduce the PDE residual instead.
[…]
where $\hat{f}=\mathcal{D}[\hat{u}_{\theta}]$ is the approximated source function. In practice, this integral is approximated via quadrature, with training error defined in equation 82.
From the theoretical perspective, the goal is to derive that the total error $\mathcal{E}_{\text{PINN}}$ equation 84 is sufficiently small. To prove this, a sufficient condition is that the total error is bounded by the residual error $\mathcal{E}_{\mathcal{R}}$ equation 100, i.e. we can prove that the smallest residual error ensures the smallest total error.
$\forall\theta\in\Theta.\;\mathcal{E}_{\text{PINN}}(\theta)\leq C\mathcal{E}_{\mathcal{R}}(\theta),$ (101)
where $C$ is a constant. By expansion of $\mathcal{E}_{\text{PINN}}$ equation 84 and $\mathcal{E}_{\mathcal{R}}$ equation 100, the abovementioned inequality equation 101 is equivalent to the following coercivity condition Ryck and Mishra (2022),
[…]
By quadrature bound Iserles (2008), the smallest practical training error $\mathcal{E}_{T}$ equation 82 ensures the smallest residual error $\mathcal{E}_{R}$ equation 100, where $C^{\prime}$ is a constant,
$\forall\theta\in\Theta.\;\mathcal{E}_{\mathcal{R}}(\theta)\leq C^{\prime}[\mathcal{E}_{T}(\theta)+\mathcal{E}_{u}(N_{\mathcal{R}})],$ (103)
and the extra term $\mathcal{E}_{u}(N_{\mathcal{R}})$ converges faster than $\frac{1}{N_{\mathcal{R}}}$, thus can be ignored given the increasing sampled quadrature points $N_{\mathcal{R}}$,
$\mathcal{E}_{u}({N_{\mathcal{R}}})\sim o(\frac{1}{{N_{\mathcal{R}}}})$ $\implies\frac{\mathcal{E}_{u}({N_{\mathcal{R}}})}{\frac{1}{{N_{\mathcal{R}}}}}=0,n\to\infty,$ $\quad\text{by the definition of little-o notation}.$ (104)
$\implies\lim_{{N_{\mathcal{R}}}\to\infty}{N_{\mathcal{R}}}\mathcal{E}_{u}({N_{\mathcal{R}}})=0,$ $\quad\text{by the definition of limit}.$
By the above two inequalities equation 101 and equation 103, the total error $\mathcal{E}_{\text{PINN}}$ equation 84 converges as the training error $\mathcal{E}_{T}$ equation 82 converges,
$\forall\theta\in\Theta.\;\mathcal{E}_{\text{PINN}}(\theta)\leq CC^{\prime}[\mathcal{E}_{T}(\theta)+o(\frac{1}{{N_{\mathcal{R}}}})].$ (105)
By Universal approximation theorem equation 95, the smoothness of the solution $u$ ensures that the residual error $\mathcal{E}_{\mathcal{R}}(\theta)<\epsilon$ is sufficiently small. Given sufficient quadrature points $N_{\mathcal{R}}$}, and smooth activation functions in the neural network $\hat{u}_{\theta}$ Iserles (2008),
$\min_{\theta\in\Theta}\mathcal{E}_{T}(\theta)\leq\mathcal{E}_{\mathcal{R}}(\theta)+o(\frac{1}{{N_{\mathcal{R}}}}),$ (106)
Hence, the training error $\mathcal{E}_{T}(\theta)<\epsilon+o(\frac{1}{{N_{\mathcal{R}}}})$ is sufficiently small, according to equation 104. So is the total error $\mathcal{E}_{\text{PINN}}(\theta)<CC^{\prime}[\epsilon+o(\frac{1}{{N_{\mathcal{R}}}})]$, by equation 105, which concludes the proof.
From the practical perspective, the common failure modes, from the above theoretical analysis, are (1) few quadrature points $N_{\mathcal{R}}$} leading to large training error $\mathcal{E}_{T}$ in equation 98, (2) insufficient training resulting in large optimization error $\mathcal{E}_{O}$ in equation 99, (3) violation of the coercivity condition equation 102 for PDEs, and (4) large constant $C$ in equation 101 or $C^{\prime}$ in equation 103.